# A general framework of Riemannian adaptive optimization methods with a convergence analysis

### v1

@@ 2.2 Riemannian stochastic optimization problem
We focus on minimizing a objective function $f:M\to\mathbb{R}$ of the form,
@@
where $f_{i}$ is a smooth function for $i=1,\ldots,N$. We use the mini-batch strategy as follows (see Iiduka (2024) for detail). $s_{k,i}$ is a random variable generated from the $i$-th sampling at the $k$-th iteration, and $\boldsymbol{s}_{k}:=(s_{k,1},\ldots,s_{k,b})^{\top}$ is independent of $(x_{k})_{k=1}^{\infty}$, where $b$ $(\leq N)$ is the batch size. To simplify the notation, we denote the expectation $\mathbb{E}_{\boldsymbol{s}_{k}}$ with respect to $\boldsymbol{s}_{k}$ by $\mathbb{E}_{k}$. From the independence of $\boldsymbol{s}_{1},\boldsymbol{s}_{2},\ldots,\boldsymbol{s}_{k}$, we can define the total expectation $\mathbb{E}$ by $\mathbb{E}_{1}\mathbb{E}_{2}\cdots\mathbb{E}_{k}$. We define the mini-batch stochastic gradient $\operatorname*{grad}f_{B_{k}}(x_{k})$ of $f$ at the $k$-th iteration by
$\operatorname*{grad}f_{B_{k}}(x_{k}):=\frac{1}{b}\sum_{i=1}^{b}\operatorname*{grad}f_{s_{k,i}}(x_{k}).$ (1)
Our main objective is to find a local minimizer of $f$, i.e., a stationary point $x_{\star}\in M$ satisfying $\operatorname*{grad}f(x_{\star})=0_{x_{\star}}$.
@@ 2.3 Proposed general framework of Riemannian adaptive methods
Although Algorithm 1 is an optimization method on Riemannian manifold $M$, since $g_{k}\in T_{x_{k}}M\subset\mathbb{R}^{d}$, $m_{k}\in\mathbb{R}^{d}$ and $H_{k}\in\mathcal{D}^{d}\cap\mathcal{S}^{d}_{++}\subset\mathbb{R}^{d\times d}$, we can directly use $(\phi_{n})_{n=1}^{\infty}$ and $(\psi_{n})_{n=1}^{\infty}$ to extend the Euclidean adaptive gradient methods.
@@
where $v_{0}=0\in\mathbb{R}^{d}$ and $\epsilon>0$. Here, we will denote the $i$-th component of $v_{k}$ by $v_{k,i}$. The exponential moving average variant of AdaGrad is often used in deep-learning training. The most basic variant is RMSProp (Hinton et al., 2012), which uses the sequences of maps $\phi_{k}(g_{1},\ldots,g_{k})=g_{k}$ and
@@
1:Initial point $x_{1}\in M$, retraction $R:TM\to M$, step sizes $(\alpha_{k})_{k=1}^{\infty}\subset\mathbb{R}_{++}$, hyperparameters $\beta_{1},\beta_{2}\in[0,1)$, $\epsilon>0$.
@@ 3.1 Assumptions and useful lemmas
    $\mathbb{E}_{k}[\operatorname*{grad}f_{s_{k,i}}(x_{k})]=\operatorname*{grad}f(x_{k})$ for all $k\geq 1$ and $i=1,\ldots,b$.
@@
Lemma 3.2.
for all $k\geq 1$.
Proof.
See Appendix A. ∎
Lemma 3.3.
$\mathbb{E}_{k}\left[\left\lVert{\operatorname*{grad}f_{B_{k}}(x_{k})}\right\rVert^{2}_{2}\right]\leq\frac{\sigma^{2}}{b}+\left\lVert{\operatorname*{grad}f(x_{k})}\right\rVert^{2}_{2}$
for all $k\geq 1$.
Proof.
See Appendix B. ∎
It is known that if Assumption 3.1 (A4) holds, so does the following Proposition 3.4. This property is known as retraction $L$-smooth (see Huang et al. (2015); Kasai et al. (2018) for details).
@@ 3.2 Convergence analysis of Algorithm 1
However, this strategy does not work in the Riemannian setting. Therefore, by following the policy of Zaheer et al. (2018), let us analyze the case in which $\phi_{k}(g_{1},\ldots,g_{k})=g_{k}$. To simplify the notation, we denote the $i$-th component of $g_{k}$ (resp. $v_{k}$, $\hat{v}_{k}$) by $g_{k,i}$ (resp. $v_{k,i}$, $\hat{v}_{k,i}$).
Lemma 3.5.
$\hat{v}_{k,i}\leq B^{2},$
for all $k\geq 1$ and $i=1,\ldots,d$.
Proof.
See Appendix C. ∎
@@
$\sum_{k=1}^{K}\alpha_{k}\left(\mu-\frac{L\alpha_{k}\nu^{2}}{2}\right)\mathbb{E}\left[\left\lVert{\operatorname*{grad}f(x_{k})}\right\rVert_{2}^{2}\right]\leq C_{1}+\frac{C_{2}}{b}\sum_{k=1}^{K}\alpha_{k}^{2},$
@@
Remark: Since Algorithm 2 satisfies $\hat{v}_{k+1,i}:=\max(\hat{v}_{k,i},v_{k+1,i})\geq\hat{v}_{k,i}$, it together with $\alpha_{k+1}\leq\alpha_{k}$, leads to $\alpha_{k}/(\sqrt{\hat{v}_{k,i}}+\epsilon)\geq\alpha_{k+1}/(\sqrt{\hat{v}_{k+1,i}}+\epsilon)$. Moreover, from Lemma 3.5,
@@
$\leq f(x_{1})+\left\langle{g(x_{1}),-\alpha_{1}\mathbb{E}_{1}[H_{1}^{-1}g_{1}]}\right\rangle_{2}+\frac{L\alpha_{1}^{2}\nu^{2}}{2}\left(\frac{\sigma^{2}}{b}+\left\lVert{g(x_{1})}\right\rVert_{2}^{2}\right)$
where the second inequality comes from Lemma 3.3. By taking $\mathbb{E}[\cdot]$ of both sides and rearranging terms, we get
$-\frac{L\alpha_{1}\nu^{2}}{2}\mathbb{E}\left[\left\lVert{g(x_{1})}\right\rVert_{2}^{2}\right]\leq f(x_{1})-\mathbb{E}[f(x_{2})]+\left\langle{g(x_{1}),-\alpha_{1}\mathbb{E}[H_{1}^{-1}g_{1}]}\right\rangle_{2}+\frac{L\alpha_{1}^{2}\sigma^{2}\nu^{2}}{2b}.$
@@
$\alpha_{1}\mu G^{2}-\frac{L\alpha_{1}\nu^{2}}{2}\mathbb{E}\left[\left\lVert{g(x_{1})}\right\rVert_{2}^{2}\right]\leq f(x_{1})-\mathbb{E}[f(x_{2})]+\frac{L\alpha_{1}^{2}\sigma^{2}\nu^{2}}{2b}+\underbrace{\left\langle{g(x_{1}),-\alpha_{1}\mathbb{E}[H_{1}^{-1}g_{1}]}\right\rangle_{2}+\alpha_{1}\mu G^{2}}_{C_{0}}.$
@@
$\alpha_{1}\left(\mu-\frac{L\alpha_{1}\nu^{2}}{2}\right)\mathbb{E}\left[\left\lVert{g(x_{1})}\right\rVert_{2}^{2}\right]\leq f(x_{1})-\mathbb{E}[f(x_{2})]+\frac{L\alpha_{1}^{2}\sigma^{2}\nu^{2}}{2b}+C_{0}.$ (4)
@@
for all $k\geq 2$. From Assumption 3.1 (A3), Lemma E.1, and $\alpha_{k-1}H_{k-1}^{-1}-\alpha_{k}H_{k}^{-1}\succeq O$, we have
@@
$\quad+GB\mathbb{E}_{k}[\operatorname*{tr}(\alpha_{k-1}H_{k-1}^{-1}-\alpha_{k}H_{k}^{-1})]+\frac{L\alpha_{k}^{2}\nu^{2}}{2}\left(\frac{\sigma^{2}}{b}+\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right),$
where the first inequality comes from the independence of $H_{k-1}^{-1}$ for $\boldsymbol{s}_{k}$ and the second inequality comes from Lemmas 3.2 and 3.3. Here, since $H_{k-1}^{-1}\succeq\mu I_{d}$ and $\alpha_{k}\leq\alpha_{k-1}$, it follows that
@@
$\mathbb{E}_{k}[f(x_{k+1})]\leq f(x_{k})-\alpha_{k}\left(\mu-\frac{L\alpha_{k}\nu^{2}}{2}\right)\left\lVert{g(x_{k})}\right\rVert_{2}^{2}+GB\mathbb{E}_{k}[\operatorname*{tr}(\alpha_{k-1}H_{k-1}^{-1}-\alpha_{k}H_{k}^{-1})]+\frac{L\alpha_{k}^{2}\sigma^{2}\nu^{2}}{2b}.$
@@
$\mathbb{E}[f(x_{k+1})]\leq\mathbb{E}[f(x_{k})]-\alpha_{k}\left(\mu-\frac{L\alpha_{k}\nu^{2}}{2}\right)\mathbb{E}\left[\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right]+GB\mathbb{E}[\operatorname*{tr}(\alpha_{k-1}H_{k-1}^{-1}-\alpha_{k}H_{k}^{-1})]+\frac{L\alpha_{k}^{2}\sigma^{2}\nu^{2}}{2b}.$
By rearranging the above inequality gives us
$\alpha_{k}\left(\mu-\frac{L\alpha_{k}\nu^{2}}{2}\right)\mathbb{E}\left[\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right]\leq\mathbb{E}[f(x_{k})]-\mathbb{E}[f(x_{k+1})]+GB\mathbb{E}[\operatorname*{tr}(\alpha_{k-1}H_{k-1}^{-1}-\alpha_{k}H_{k}^{-1})]+\frac{L\alpha_{k}^{2}\sigma^{2}\nu^{2}}{2b}.$ (5)
@@
$\quad\leq\mathbb{E}[f(x_{2})]-\mathbb{E}[f(x_{K+1})]+GB\mathbb{E}[\operatorname*{tr}(\alpha_{1}H_{1}^{-1}-\alpha_{K}H_{K}^{-1})]+\sum_{k=2}^{K}\frac{L\alpha_{k}^{2}\sigma^{2}\nu^{2}}{2b}.$
@@
$\sum_{k=2}^{K}\alpha_{k}\left(\mu-\frac{L\alpha_{k}\nu^{2}}{2}\right)\mathbb{E}\left[\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right]\leq\mathbb{E}[f(x_{2})]-f_{\star}+GB\alpha_{1}\nu d+\sum_{k=2}^{K}\frac{L\alpha_{k}^{2}\sigma^{2}\nu^{2}}{2b}.$ (6)
@@
$\sum_{k=1}^{K}\alpha_{k}\left(\mu-\frac{L\alpha_{k}\nu^{2}}{2}\right)\mathbb{E}\left[\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right]\leq\underbrace{f(x_{1})-f_{\star}+C_{0}+GB\alpha_{1}\nu d}_{C_{1}}+\underbrace{\frac{L\sigma^{2}\nu^{2}}{2}}_{C_{2}}\cdot\frac{1}{b}\sum_{k=1}^{K}\alpha_{k}^{2}.$
@@
Our convergence analysis (Theorem 3.7) allows the proposed framework (Algorithm 1) to use both constant and diminishing steps sizes. Theorems 3.8 and 3.9 are convergence analyses of Algorithm 1 with constant and diminishing steps sizes, respectively.
@@
Under the assumptions in Theorem 3.7 and assuming that the constant step size $\alpha_{k}:=\alpha$ satisfies $0<\alpha<2\mu L^{-1}\nu^{-2}$, the sequence $(x_{k})_{k=1}^{\infty}\subset M$ generated by Algorithm 1 satisfies
@@
Under the assumptions in Theorem 3.7 and assuming that the diminishing step size $\alpha_{k}:=\alpha/\sqrt{k}$ satisfies $\alpha\in(0,1]$, the sequence $(x_{k})_{k=1}^{\infty}\subset M$ generated by Algorithm 1 satisfies

### v2

@@ 2.2 Riemannian stochastic optimization problem
We focus on minimizing an objective function $f:M\to\mathbb{R}$ of the form,
@@
where $f_{i}$ is a smooth function for $i=1,\ldots,N$. We use the mini-batch strategy as follows (see Iiduka (2024) for detail). $s_{k,i}$ is a random variable generated from the $i$-th sampling at the $k$-th iteration, and $\boldsymbol{s}_{k}:=(s_{k,1},\ldots,s_{k,b_{k}})^{\top}$ is independent of $(x_{k})_{k=1}^{\infty}$, where $b_{k}$ $(\leq N)$ is the batch size. To simplify the notation, we denote the expectation $\mathbb{E}_{\boldsymbol{s}_{k}}$ with respect to $\boldsymbol{s}_{k}$ by $\mathbb{E}_{k}$. From the independence of $\boldsymbol{s}_{1},\boldsymbol{s}_{2},\ldots,\boldsymbol{s}_{k}$, we can define the total expectation $\mathbb{E}$ by $\mathbb{E}_{1}\mathbb{E}_{2}\cdots\mathbb{E}_{k}$. We define the mini-batch stochastic gradient $\operatorname*{grad}f_{B_{k}}(x_{k})$ of $f$ at the $k$-th iteration by
$\operatorname*{grad}f_{B_{k}}(x_{k}):=\frac{1}{b_{k}}\sum_{i=1}^{b_{k}}\operatorname*{grad}f_{s_{k,i}}(x_{k}).$ (1)
Our main objective is to find a stationary point $x_{\star}\in M$ satisfying $\operatorname*{grad}f(x_{\star})=0_{x_{\star}}$.
@@ 2.3 Proposed general framework of Riemannian adaptive methods
By setting the maps $(\phi_{n})_{n=1}^{\infty}$ and $(\psi_{n})_{n=1}^{\infty}$ in Algorithm 1 to those used by an adaptive optimization method in Euclidean space, any adaptive optimization method can be extended to Riemannian manifolds. In the following, we show some examples of extending the optimization algorithm in Euclidean space to Riemannian manifolds by using Algorithm 1.
@@
where $v_{0}=0\in\mathbb{R}^{d}$ and $\epsilon>0$. Here, we will denote the $i$-th component of $v_{k}$ by $v_{k,i}$. The exponential moving average variant of AdaGrad is often used in deep-learning training. The basic variant is RMSProp (Hinton et al., 2012), which uses the sequences of maps $\phi_{k}(g_{1},\ldots,g_{k})=g_{k}$ and
@@
1:Initial point $x_{1}\in M$, retraction $R:TM\to M$, step sizes $(\alpha_{k})_{k=1}^{\infty}\subset\mathbb{R}_{++}$, mini-batch sizes $(b_{k})_{k=1}^{\infty}\subset\mathbb{R}_{++}$ hyperparameters $\beta_{1},\beta_{2}\in[0,1)$, $\epsilon>0$.
@@ 3.1 Assumptions and useful lemmas
    $\mathbb{E}_{k}[\operatorname*{grad}f_{s_{k,i}}(x_{k})]=\operatorname*{grad}f(x_{k})$ for all $k\geq 1$ and $i=1,\ldots,b_{k}$.
@@
Assumptions (A1) and (A2) are satisfied when the probability distribution is uniform (see Appendix F) or when the full gradient is used, i.e., when the batch size equals the total number of data. (A1) and (A2) have been used to analyze adaptive methods in Euclidean space (Zaheer et al., 2018, Pages 3, 12, and 16) and on the Hadamard manifold (Sakai & Iiduka, 2024, (2.2) and (2.3)). Assumptions (A3)(A5) are often used in both Euclidean spaces and Riemannian manifolds. Intuitively, in practical applications, these assumptions are satisfied because the search space of the algorithm is bounded. If $f$ is unbounded on an unbounded search space $X$, then (A5) does not hold. Although there is a possibility such that (A4) is satisfied, (A2) would not hold, even when the probability distribution is uniform. When $f$ is unbounded, there is no minimizer of $f$. Hence, (A5) is essential in order to analyze optimization methods. Moreover, (A4) is needed in order to use a retraction $L$-smooth (Proposition 3.2) that is an extension of the Euclidean-type descent lemma. (A3) holds when the manifold is compact (Absil et al., 2008) or through a slight modification of the objective function and the algorithm (Kasai et al., 2018).
It is known that if Assumption 3.1 (A4) holds, so does the following Proposition 3.2. This property is known as retraction $L$-smoothness (see Huang et al. (2015); Kasai et al. (2018) for details).
@@ 3.2 Convergence analysis of Algorithm 1
However, since the embedded submanifold is not closed under addition in Euclidean space, this strategy does not work in the Riemannian setting. Therefore, by following the policy of Zaheer et al. (2018), let us analyze the case in which $\phi_{k}(g_{1},\ldots,g_{k})=g_{k}$. To simplify the notation, we denote the $i$-th component of $g_{k}$ (resp. $v_{k}$, $\hat{v}_{k}$) by $g_{k,i}$ (resp. $v_{k,i}$, $\hat{v}_{k,i}$).
@@
$\sum_{k=1}^{K}\alpha_{k}\left(\mu-\frac{L\alpha_{k}\nu^{2}}{2}\right)\mathbb{E}\left[\left\lVert{\operatorname*{grad}f(x_{k})}\right\rVert_{2}^{2}\right]\leq C_{1}+C_{2}\sum_{k=1}^{K}\frac{\alpha_{k}^{2}}{b_{k}},$
@@
Remark: Since Algorithm 2 satisfies $\hat{v}_{k+1,i}:=\max(\hat{v}_{k,i},v_{k+1,i})\geq\hat{v}_{k,i}$, it together with $\alpha_{k+1}\leq\alpha_{k}$, leads to $\alpha_{k}/(\sqrt{\hat{v}_{k,i}}+\epsilon)\geq\alpha_{k+1}/(\sqrt{\hat{v}_{k+1,i}}+\epsilon)$. Moreover, from Lemma A.3,
@@
$\leq f(x_{1})+\left\langle{g(x_{1}),-\alpha_{1}\mathbb{E}_{1}[H_{1}^{-1}g_{1}]}\right\rangle_{2}+\frac{L\alpha_{1}^{2}\nu^{2}}{2}\left(\frac{\sigma^{2}}{b_{1}}+\left\lVert{g(x_{1})}\right\rVert_{2}^{2}\right)$
where the second inequality comes from Lemma A.2. By taking $\mathbb{E}[\cdot]$ of both sides and rearranging terms, we get
$-\frac{L\alpha_{1}\nu^{2}}{2}\mathbb{E}\left[\left\lVert{g(x_{1})}\right\rVert_{2}^{2}\right]\leq f(x_{1})-\mathbb{E}[f(x_{2})]+\left\langle{g(x_{1}),-\alpha_{1}\mathbb{E}[H_{1}^{-1}g_{1}]}\right\rangle_{2}+\frac{L\alpha_{1}^{2}\sigma^{2}\nu^{2}}{2b_{1}}.$
@@
$\alpha_{1}\mu G^{2}-\frac{L\alpha_{1}\nu^{2}}{2}\mathbb{E}\left[\left\lVert{g(x_{1})}\right\rVert_{2}^{2}\right]\leq f(x_{1})-\mathbb{E}[f(x_{2})]+\frac{L\alpha_{1}^{2}\sigma^{2}\nu^{2}}{2b_{1}}+\underbrace{\left\langle{g(x_{1}),-\alpha_{1}\mathbb{E}[H_{1}^{-1}g_{1}]}\right\rangle_{2}+\alpha_{1}\mu G^{2}}_{C_{0}}.$
@@
$\alpha_{1}\left(\mu-\frac{L\alpha_{1}\nu^{2}}{2}\right)\mathbb{E}\left[\left\lVert{g(x_{1})}\right\rVert_{2}^{2}\right]\leq f(x_{1})-\mathbb{E}[f(x_{2})]+\frac{L\alpha_{1}^{2}\sigma^{2}\nu^{2}}{2b_{1}}+C_{0}.$ (4)
@@
for all $k\geq 2$. From Assumption 3.1 (A3), Lemma C.1, and $\alpha_{k-1}H_{k-1}^{-1}-\alpha_{k}H_{k}^{-1}\succeq O$, we have
@@
$\quad+GB\mathbb{E}_{k}[\operatorname*{tr}(\alpha_{k-1}H_{k-1}^{-1}-\alpha_{k}H_{k}^{-1})]+\frac{L\alpha_{k}^{2}\nu^{2}}{2}\left(\frac{\sigma^{2}}{b_{k}}+\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right),$
where the first inequality comes from the independence of $H_{k-1}^{-1}$ for $\boldsymbol{s}_{k}$ and the second inequality comes from Lemmas A.1 and A.2. Here, since $H_{k-1}^{-1}\succeq\mu I_{d}$ and $\alpha_{k}\leq\alpha_{k-1}$, it follows that
@@
$\mathbb{E}_{k}[f(x_{k+1})]\leq f(x_{k})-\alpha_{k}\left(\mu-\frac{L\alpha_{k}\nu^{2}}{2}\right)\left\lVert{g(x_{k})}\right\rVert_{2}^{2}+GB\mathbb{E}_{k}[\operatorname*{tr}(\alpha_{k-1}H_{k-1}^{-1}-\alpha_{k}H_{k}^{-1})]+\frac{L\alpha_{k}^{2}\sigma^{2}\nu^{2}}{2b_{k}}.$
@@
$\mathbb{E}[f(x_{k+1})]\leq\mathbb{E}[f(x_{k})]-\alpha_{k}\left(\mu-\frac{L\alpha_{k}\nu^{2}}{2}\right)\mathbb{E}\left[\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right]+GB\mathbb{E}[\operatorname*{tr}(\alpha_{k-1}H_{k-1}^{-1}-\alpha_{k}H_{k}^{-1})]+\frac{L\alpha_{k}^{2}\sigma^{2}\nu^{2}}{2b_{k}}.$
Rearranging the above inequality gives us
$\alpha_{k}\left(\mu-\frac{L\alpha_{k}\nu^{2}}{2}\right)\mathbb{E}\left[\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right]\leq\mathbb{E}[f(x_{k})]-\mathbb{E}[f(x_{k+1})]+GB\mathbb{E}[\operatorname*{tr}(\alpha_{k-1}H_{k-1}^{-1}-\alpha_{k}H_{k}^{-1})]+\frac{L\alpha_{k}^{2}\sigma^{2}\nu^{2}}{2b_{k}}.$ (5)
@@
$\quad\leq\mathbb{E}[f(x_{2})]-\mathbb{E}[f(x_{K+1})]+GB\mathbb{E}[\operatorname*{tr}(\alpha_{1}H_{1}^{-1}-\alpha_{K}H_{K}^{-1})]+\sum_{k=2}^{K}\frac{L\alpha_{k}^{2}\sigma^{2}\nu^{2}}{2b_{k}}.$
@@
$\sum_{k=2}^{K}\alpha_{k}\left(\mu-\frac{L\alpha_{k}\nu^{2}}{2}\right)\mathbb{E}\left[\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right]\leq\mathbb{E}[f(x_{2})]-f_{\star}+GB\alpha_{1}\nu d+\sum_{k=2}^{K}\frac{L\alpha_{k}^{2}\sigma^{2}\nu^{2}}{2b_{k}}.$ (6)
@@
$\sum_{k=1}^{K}\alpha_{k}\left(\mu-\frac{L\alpha_{k}\nu^{2}}{2}\right)\mathbb{E}\left[\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right]\leq\underbrace{f(x_{1})-f_{\star}+C_{0}+GB\alpha_{1}\nu d}_{C_{1}}+\underbrace{\frac{L\sigma^{2}\nu^{2}}{2}}_{C_{2}}\sum_{k=1}^{K}\frac{\alpha_{k}^{2}}{b_{k}}.$
@@
Our convergence analysis (Theorem 3.4) allows the proposed framework (Algorithm 1) to use both constant and diminishing steps sizes. Theorems 3.5 and 3.6 are convergence analyses of Algorithm 1 with constant and diminishing steps sizes and a fixed mini-batch size, respectively.
@@
Under the assumptions in Theorem 3.4 and assuming that the constant step size $\alpha_{k}:=\alpha$ satisfies $0<\alpha<2\mu L^{-1}\nu^{-2}$ and the mini-batch size $b_{k}:=b$ satisfies $0<b\leq N$, the sequence $(x_{k})_{k=1}^{\infty}\subset M$ generated by Algorithm 1 satisfies
@@
Under the assumptions in Theorem 3.4 and assuming that the diminishing step size $\alpha_{k}:=\alpha/\sqrt{k}$ satisfies $\alpha\in(0,1]$ and the mini-batch size $b_{k}:=b$ satisfies $0<b\leq N$, the sequence $(x_{k})_{k=1}^{\infty}\subset M$ generated by Algorithm 1 satisfies
@@
Theorems 3.7 and 3.8 are convergence analyses of Algorithm 1 with constant and diminishing steps sizes and increasing mini-batch sizes, respectively.
Theorem 3.7.
Under the assumptions in Theorem 3.4 and assuming that the constant step size $\alpha_{k}:=\alpha$ satisfies $0<\alpha<2\mu L^{-1}\nu^{-2}$ and the mini-batch size $b_{k}$ satisfies $0<b_{k}\leq N$ and $\sum_{k=1}^{\infty}b_{k}^{-1}<\infty$, the sequence $(x_{k})_{k=1}^{\infty}\subset M$ generated by Algorithm 1 satisfies
$\min_{k=1,\ldots,K}\mathbb{E}\left[\left\lVert{\operatorname*{grad}f(x_{k})}\right\rVert_{2}^{2}\right]\leq\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\left[\left\lVert{\operatorname*{grad}f(x_{k})}\right\rVert_{2}^{2}\right]=\mathcal{O}\left(\frac{1}{K}\right).$
Proof.
$\frac{1}{K}\sum_{k=1}^{K}\alpha\left(\mu-\frac{L\alpha\nu^{2}}{2}\right)\mathbb{E}\left[\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right]\leq\frac{C_{1}}{K}+\frac{C_{2}\alpha^{2}}{K}\sum_{k=1}^{K}\frac{1}{b_{k}}.$ (8)
$\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\left[\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right]\leq\frac{2C_{1}}{2\alpha\mu-L\alpha^{2}\nu^{2}}\cdot\frac{1}{K}+\frac{2C_{2}\alpha^{2}}{2\alpha\mu-L\alpha^{2}\nu^{2}}\cdot\frac{1}{K}\sum_{k=1}^{K}\frac{1}{b_{k}}.$
This completes the proof. ∎
An example of $b_{k}$ such that $\sum_{k=1}^{\infty}b_{k}^{-1}<\infty$ used in Theorem 3.7 is such that the mini-batch size is multiplied by $\delta>1$ every step, i.e.,
$b_{t}=\delta^{t}b_{0}.$ (9)
For example, the mini-batch size defined by (9) with $\delta=2$ makes batch size double each step. Moreover, the mini-batch size defined by (9) and the diminishing step size $\alpha_{k}:=\alpha/\sqrt{k}$, where $\alpha\in(0,1]$, satisfy that $\sum_{k=1}^{\infty}\alpha_{k}^{2}b_{k}^{-1}<\infty$ (see the assumption in Theorem 3.8).
Theorem 3.8.
Under the assumptions in Theorem 3.4 and assuming that the diminishing step size $\alpha_{k}:=\alpha/\sqrt{k}$ satisfies $\alpha\in(0,1]$ and the mini-batch size $b_{k}$ satisfies $0<b_{k}\leq N$ and $\sum_{k=1}^{\infty}\alpha_{k}^{2}b_{k}^{-1}<\infty$, the sequence $(x_{k})_{k=1}^{\infty}\subset M$ generated by Algorithm 1 satisfies
$\min_{k=1,\ldots,K}\mathbb{E}\left[\left\lVert{\operatorname*{grad}f(x_{k})}\right\rVert_{2}^{2}\right]\leq\frac{1}{\sum_{k=1}^{K}\alpha_{k}}\sum_{k=1}^{K}\alpha_{k}\mathbb{E}\left[\left\lVert{\operatorname*{grad}f(x_{k})}\right\rVert_{2}^{2}\right]=\mathcal{O}\left(\frac{1}{\sqrt{K}}\right).$
Proof.
$0<\mu-\frac{L\alpha_{k}\nu^{2}}{2}<\mu,$
for all $k\geq k_{0}$. From Theorem 3.4, we have
$\sum_{k=k_{0}}^{K}\alpha_{k}\left(\mu-\frac{L\alpha_{k}\nu^{2}}{2}\right)\mathbb{E}\left[\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right]\leq C_{1}+C_{2}\sum_{k=1}^{K}\frac{\alpha_{k}^{2}}{b_{k}}-\sum_{k=1}^{k_{0}-1}\alpha_{k}\left(\mu-\frac{L\alpha_{k}\nu^{2}}{2}\right)\mathbb{E}\left[\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right],$
$\left(\mu-\frac{L\alpha_{k_{0}}\nu^{2}}{2}\right)\sum_{k=k_{0}}^{K}\alpha_{k}\mathbb{E}\left[\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right]\leq C_{1}+C_{2}\sum_{k=1}^{K}\frac{\alpha_{k}^{2}}{b_{k}}+\sum_{k=1}^{k_{0}-1}L\alpha_{k}^{2}\nu^{2}\mathbb{E}\left[\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right].$
Dividing both sides of this inequality by $02^{-1}(2\mu-L\alpha_{k_{0}}\nu^{2})>$ 0 yields
$\sum_{k=k_{0}}^{K}\alpha_{k}\mathbb{E}\left[\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right]$ $\leq\frac{2}{2\mu-L\alpha_{k_{0}}\nu^{2}}\left(C_{1}+C_{2}\sum_{k=1}^{K}\frac{\alpha_{k}^{2}}{b_{k}}+\sum_{k=1}^{k_{0}-1}L\alpha_{k}^{2}\nu^{2}\mathbb{E}\left[\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right]\right)$
$=\underbrace{\frac{2}{2\mu-L\alpha_{k_{0}}\nu^{2}}\left(C_{1}+\sum_{k=1}^{k_{0}-1}L\alpha_{k}^{2}\nu^{2}\mathbb{E}\left[\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right]\right)}_{C_{3}}$
$\quad+\underbrace{\frac{2C_{2}}{2\mu-L\alpha_{k_{0}}\nu^{2}}}_{C_{4}}\sum_{k=1}^{K}\frac{\alpha_{k}^{2}}{b_{k}}.$
From this and $\sum_{k=1}^{\infty}\alpha_{k}^{2}{b_{k}}^{-1}<\infty$, we have that
$\sum_{k=1}^{K}\alpha_{k}\mathbb{E}\left[\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right]$ $\leq\underbrace{C_{3}+\sum_{k=1}^{k_{0}-1}\alpha_{k}\mathbb{E}\left[\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right]+C_{4}\sum_{k=1}^{K}\frac{\alpha_{k}^{2}}{b_{k}}}_{C_{5}},$
which implies that
$\frac{1}{\sum_{k=1}^{K}\alpha_{k}}\sum_{k=1}^{K}\alpha_{k}\mathbb{E}\left[\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right]$ $\leq\frac{C_{5}}{\sum_{k=1}^{K}\alpha_{k}}.$
From $\alpha_{k}:=\alpha/\sqrt{k}$, we obtain
$\sum_{k=1}^{K}\alpha_{k}=\sum_{k=1}^{K}\frac{\alpha}{\sqrt{k}}\geq\alpha\int_{1}^{K+1}\frac{dt}{\sqrt{t}}=2\alpha(\sqrt{K+1}-1).$
Therefore,
$\frac{1}{\sum_{k=1}^{K}\alpha_{k}}\sum_{k=1}^{K}\alpha_{k}\mathbb{E}\left[\left\lVert{g(x_{k})}\right\rVert_{2}^{2}\right]\leq\frac{C_{5}}{2\alpha(\sqrt{K+1}-1)}.$
This completes the proof. ∎